OpenStudy (anonymous):

Pweeez help me): Please help me ! Solve x^2 - 3x = -8. x equals quantity of 3 plus or minus I square root of 29 all over 2 x equals quantity of 3 plus or minus I square root of 23 all over 2 x equals quantity of negative 3 plus or minus I square root of 29 all over 2 x equals quantity of negative 3 plus or minus I square root of 23 all over 2

OpenStudy (michelleceirra606):

-1x2 + x2 = 0 -8 + -3x + x2 = 0